OpenStudy (anonymous):

Write an equation in y = mx + b form for the table. # books purchased 1 2 3 4 Cost of book club 21 30 39 48 A. y = 3x + 21 B. y = 9x + 12 C. y = 5x + 21 D. y = 7x + 12

OpenStudy (anonymous):

m=(y'-y)/(x'-x) =(30-21)/(2-1) =9 y=9x+b 21=9*1+b 12=b

OpenStudy (anonymous):

plug in and see....

OpenStudy (anonymous):

b.